OpenStudy (kayders1997):

Let a function g be defined by g(x)=2x^2-3x+10 find g'(x) when g(x)=15

OpenStudy (kayders1997):

So 2x^2-3x+10=15? Solve for x?

OpenStudy (caozeyuan):

Yes

OpenStudy (kayders1997):

I'm just confused thought because you get more than 1 point? Is that ok?

OpenStudy (caozeyuan):

Of ccause, and their sloe should be direct opposite

OpenStudy (caozeyuan):

sorry, slope, my typo

OpenStudy (caozeyuan):

becuase you are looking at a quartic equation

OpenStudy (caozeyuan):



OpenStudy (kayders1997):

Ok let me try it out and see what I get

OpenStudy (caozeyuan):

baddrawing but you get the idea

OpenStudy (kayders1997):

Ok so after factoring I got x=5/2 and x=-1

OpenStudy (zzr0ck3r):

plug those into the derivative

OpenStudy (kayders1997):

So I got g'(5/2)=7 and g'(1)=-7

OpenStudy (caozeyuan):

see, thats exactly what Im talking about

OpenStudy (kayders1997):

Yes, that totally made sense when you interpreted it with a graph

OpenStudy (kayders1997):

Thank you so much :)

OpenStudy (caozeyuan):

no problem, now Im going to sleep before my head hits my keyboards;dl'fmceiksd'icmmkal'sklaskdnrfsow'kqd012 e][;sdlfrpgerf[4

OpenStudy (kayders1997):

Lol goodnoght!